OpenStudy (anonymous):

Factor x^3+8x^2+19x+12. I got (x+4)(x+2)(x+6) am i right?

OpenStudy (anonymous):

nope check again...

OpenStudy (anonymous):

aagaye....aaj der lagadi

OpenStudy (anonymous):

okay I got (x+2)(x+6)(x+12)...

OpenStudy (anonymous):

I don't think that is completely right...

OpenStudy (anonymous):

(x+4)(x+3)(x+1)

OpenStudy (anonymous):

Your (x+4) was right I normally use the remainder theorum to sub in values So to check (x+4) is a factor you use x=-4 (the solution to x+4=0 into x^3+8x^2+19x+12 and it'll equal 0 if it is a factor Then use long division to leave you with a quadratic (x+4)(x^2+4x+3) then factories the quadratic